The average grade/mark of all college students is 70 with standard deviation of 20. If a random sample of 49 students is taken, what is the probability that the sample average grade/mark is
a) At least 65
b) Less than 78
c) Between 65 and 72
1
Expert's answer
2020-05-19T18:55:45-0400

"Given\;that\, \bar x=70,s=20,n=49\\
(a) P(\bar x>65)=P(Z>\frac{65-70}{\frac{20}{\sqrt{49}}})\\
=P(Z>-1.75)=0.5+P(0<Z<1.75)\\
=0.5+0.4599=0.9599\\
(b) P(\bar x<78)=P(Z<\frac{78-70}{\frac{20}{\sqrt{49}}})\\
=P(Z<2.8)=0.5+P(0<Z<2.8)\\
=0.5+0.4974=0.9974\\
 (c)P(65<\bar x<72)=P(\frac{65-70}{\frac{20}{\sqrt{49}}})<Z<\frac{72-70}{\frac{20}{\sqrt{49}}})\\
=P(-1.75<Z<0.7)\\
=P(0<Z<1.75)+P(0<Z<0.7)\\
=0.4599+0.2580=0.7179\\"